OpenStudy (anonymous):

4log3 32 (first 3 is subscript) divided by 5 log3 4 (3 is base)

OpenStudy (ash2326):

\[\frac{4\log_3 32}{5\log_3 4}\] Is this your question?

OpenStudy (anonymous):

is the answer 2

OpenStudy (anonymous):

@ash2326 , yes it is my question

OpenStudy (ash2326):

So you know the basic properties of logarithms ? @teeny

OpenStudy (callisto):

Hints <1> \(32 = 2^5\) and \(4=2^2\) <2> \(log_ab^c = c\log_ab\)